OpenStudy (anonymous):

1. Calculate the discriminant. x2 – 2x + 4 = 0 2. What number would you have to add to both sides to complete the square? x2 – 14x = 7 3. Solve by using the perfect squares method. x2 – 12x + 36 = 0 4. Solve x2 + x – 6 = 0 5. Atasha is solving the equation x2 – x = –3 with the quadratic formula. What values should she use for a, b, and c? 6. Solve (8x + 3)2 = 16 7. Use the discriminant to determine the number and type of solutions the equation has. x2 + 8x + 12 = 0

OpenStudy (blank ):

1. -12 2. 49 3. x = 6 4. x=-3,2 5. a=1, b=−1, c=3 6. x=1/8, -7/8 7. x=-6,-2